### v1

# IndoCulture: Exploring Geographically-Influenced Cultural Commonsense Reasoning Across Eleven Indonesian Provinces

Abstract
Although commonsense reasoning is greatly shaped by cultural and geographical factors, previous studies on language models have predominantly centered on English cultures, potentially resulting in an Anglocentric bias. In this paper, we introduce IndoCulture, aimed at understanding the influence of geographical factors on language model reasoning ability, with a specific emphasis on the diverse cultures found within eleven Indonesian provinces. In contrast to prior works that relied on template Yin et al. (2022) and online scrapping Fung et al. (2024), we create IndoCulture by asking local people to manually develop the context and plausible options based on predefined topics. Evaluations of 23 language models reveal several insights: (1) even the best open-source model struggles with an accuracy of 53.2%, (2) models often provide more accurate predictions for specific provinces, such as Bali and West Java, and (3) the inclusion of location contexts enhances performance, especially in larger models like GPT-4, emphasizing the significance of geographical context in commonsense reasoning. [^1] IndoCulture can be accessed at https://huggingface.co/datasets/indolem/IndoCulture
1 Introduction
Figure 1: IndoCulture covers eleven provinces spanning from eastern to western Indonesia. The highlighted regions in the map represent the provinces examined in IndoCulture. We offer examples from Aceh, South Borneo, and Papua, with correct answers indicated in bold. English translations are for illustrative purposes.
Reasoning abilities of multilingual language models are frequently evaluated using English texts, potentially amplifying an Anglocentric bias toward English culture and leading to less inclusive models Thomas (1983); Ponti et al. (2020). Cultures, however, vary significantly from one location to another and profoundly shape the way a speaker of a language reasons Hershcovich et al. (2022). Recent evaluations of models reasoning ability OpenAI (2023); Touvron et al. (2023); Jiang et al. (2023); Sengupta et al. (2023); Liu et al. (2023b) conducted on the English datasets, such as PIQA Bisk et al. (2020a), HellaSwag Zellers et al. (2019), Social IQA Sap et al. (2019), often overlook geographical aspects, thereby risking cultural bias.
Culture is a multifaceted concept encompassing the way of life Giddens and Sutton (2021), including our thoughts and actions Macionis John . It includes tangible elements like food art, and clothing, as well as intangible aspects such as ideas, values, attitudes, and norms. Culture is shaped by geographical location and ethnicity, influencing the commonsense reasoning of people within a region. For example, in Indonesia, it is culturally acceptable to eat rice with your hands but it is considered unusual to use a fork. Similarly, at traditional Indonesian weddings, it is common to sit on the floor while eating, whereas this practice is less common in Australia.
This work focuses on understanding the influence of geographical contexts in cultural commonsense reasoning, with the main focus on Indonesian culture. Indonesia is a highly multicultural country Putra et al. (2019), home to over 1,300 recognized ethnic groups and more than 700 languages Zarbaliyev (2017); Aji et al. (2022). As the the largest archipelagic country in the world, Indonesia has a population exceeding 270 million spread across 38 provinces, stretching from Aceh province in the west to Papua province in the east. Few prior studies on commonsense reasoning in Indonesian contexts Mahendra et al. (2021); Wibowo et al. (2023); Putri et al. (2024) overlooked the geographical nuance and the rich diversity of Indonesian cultures.
This paper introduces IndoCulture, a novel dataset to evaluate cultural reasoning in eleven Indonesian provinces, manually developed by local people in each province based on predefined topics. In prior works, cultural reasoning has primarily relied on datasets constructed through templates Yin et al. (2022), and online scraping Nguyen et al. (2023a); Fung et al. (2024). While these studies offer valuable insights, they may be susceptible to training data contamination when utilized to assess large language models (LLMs). For instance, Fung et al. (2024) reported a zero-shot accuracy of 92% when using ChatGPT to evaluate low-resource data.
Figure 1 illustrates the eleven provinces examined in IndoCulture including Aceh, North Sumatra, West Sumatra, West Java, Central Java, East Java, Bali, South Borneo, East Nusa Tenggara (NTT), South Sulawesi, and Papua. [^2] Although Papua consists of six provinces, for the purpose of this study, we treat it as a single entity (referred to as Papua) due to the relatively recent establishment of most of these provinces. These provinces span from the eastern to the western parts of Indonesia, each representing a major island in the country, with the addition of Bali and NTT. Figure 1 also shows three examples in IndoCulture for three provinces: Aceh, South Borneo, and Papua. The first two examples focus on cultural artifacts: the traditional wedding dress from Aceh and the Bagandut dance from South Borneo. The second example also examines cultural symbols. The third example focuses on cultural beliefs and norms regarding pregnancy in Papua.
Can large language models effectively reason based on the diverse cultures of Indonesia? To capture the rich diversity of Indonesian cultures, we meticulously predefined 12 fine-grained topics as guidelines for data construction. Figure 2 displays the topic distribution in IndoCulture, with the majority focusing on food, weddings, art, pregnancy and children, and family relationships. Additionally, we also pose the question: Is there an influence of geographical location on commonsense reasoning by language models? We address these questions through comprehensive experiments across different language models, incorporating several levels of location granularity as additional context in the prompt.
Our contributions can be summarized as follows:
  • We present IndoCulture, a high-quality cultural reasoning dataset in the Indonesian language, covering eleven provinces of Indonesia and twelve fine-grained cultural topics. Our dataset has 2,429 instances, and was meticulously developed by local people with rigorous quality controls in place.
  • We assess 15 open-source multilingual models, 6 open-source Indonesian-centric models, and 2 closed-source models. Although local individuals can answer all questions correctly (i.e., 100% accuracy), we observed that all open-source models struggle to comprehend Indonesian cultures, with the highest performance reaching only 53.2%.
  • We conduct a thorough analysis over various dimensions: (1) models performance for each province and topics, (2) the influence of different granularities of location contexts (i.e., none, province, country), (3) model performance over English translation, (4) analysis over model explanation behind its answer.
2 Related Work
Commonsense Reasoning in English
Many studies have focused on commonsense reasoning in English, often overlooking considerations of cultures and geographical locations. Early works include the Winograd Schema Challenge Levesque et al. (2012) and WinoGrande Sakaguchi et al. (2021) for pronoun coreference resolution tasks. Other research areas include reasoning based on cause-effect relationships Roemmele et al. (2011), physical activities Bisk et al. (2020b), social interaction Sap et al. (2019), cloze story completion Mostafazadeh et al. (2016), sentence completion Zellers et al. (2019), numerical Lin et al. (2020), and temporal Qin et al. (2021). Additionally, pretrained language models are employed in other works to extract structured commonsense knowledge by providing seed words Davison et al. (2019), and using code language model Madaan et al. (2022).
Cultural Commonsense Reasoning with Geographical Contexts
Previous studies have explored commonsense reasoning with geographical contexts. Shwartz (2022) investigated time perception (e.g., morning and night hours) across different locations, while Yin et al. (2022) delved into cultural knowledge of language models across five countries by masking and template strategies. Other works have focused on automatically extracting cultural knowledge from various sources, including Wikipedia Fung et al. (2024), conversations Fung et al. (2023), and Common Crawl Nguyen et al. (2023a), incorporating location contexts with the assistance of large language models. Related, Ziems et al. (2023) created a knowledge bank for situational norms, utilizing English-speaking Mechanical Turk annotators and incorporating country taxonomy. Unlike these works, IndoCulture specifically concentrates on cultural reasoning within Indonesian provinces, developed and validated manually by local people (experts). Compared to the automatic method and English-speaking crowdsource workers for data construction, IndoCulture arguably contains less noise, and is free from training data contamination of large language models (LLMs).
Commonsense Reasoning with Indonesian contexts
Commonsense reasoning in Indonesian language models has been studied using translated English–Indonesian datasets, such as XCOPA Ponti et al. (2020) and XStoryCloze Lin et al. (2021). However, these datasets potentially introduce a cultural bias toward English culture. IndoCloze Koto et al. (2022) is the first commonsense reasoning dataset in Indonesian, developed by native Indonesian workers following cloze story completion framework Mostafazadeh et al. (2016). IndoMMLU Koto et al. (2023), a benchmark for assessing knowledge in the Indonesian context, includes questions pertaining to Indonesian culture. However, its primary focus lies on testing proficiency in the local language rather than reasoning abilities. In a contemporary study, Wibowo et al. (2023) followed the COPA framework Roemmele et al. (2011) to build a dataset with contexts limited to Jakarta. In other work, Putri et al. (2024) studied the capability of LLMs in generating questions with cultural norms, for both general Indonesian and specific Sundanese contexts. In comparison, IndoCulture offers broader coverage across eleven provinces, thereby fostering greater inclusivity for local communities in Indonesia.
3 IndoCulture
As illustrated in Figure 1, IndoCulture is a sentence completion task in Indonesian language featuring a one-sentence premise, three plausible options, and one correct option to evaluate reasoning ability and cultural knowledge across eleven Indonesian provinces. In total, there are 2,429 instances, with the distribution across provinces detailed in Table 1.
3.1 Data Construction
IndoCulture is constructed manually by human and verified through a two-step process.
Worker Recruitment
Culture often arises from the shared experiences, traditions, and beliefs of a specific group over time, often closely intertwined with native populations. Keeping this in mind, we engaged individuals from various provinces across Indonesia to assist in preparing data for the IndoCulture benchmark. During the recruitment, we presented a few examples of the intended IndoCulture data and requested each candidate to generate similar instances tailored to the context of their respective provinces. From a pool of 58 applicants, we carefully selected 22 expert workers representing 11 provinces (with 2 workers selected per province). These recruited expert workers are local residents and have resided in their respective provinces for a minimum of 10 years, thereby possessing a profound understanding of local customs and culture. The age range of our workforce spans from 21 to 35 years old, with educational backgrounds distributed as follows: 3 high school graduates, 14 bachelor’s degree holders, 4 master’s degree holders, and 1 PhD holder. During data construction, each expert worker will fulfill the dual roles of instance writer and quality controller. [^3] Each worker is compensated above the monthly minimum wage in Indonesia.
Province Selection
The provinces covered in this study represent diversity of Indonesian cultures. Those 11 provinces (in Figure 1) spread across 6 islands and large archipelagos in Indonesia which are inhabited by various tribes who speak different regional languages and adhere to different religions.
Topic Taxonomy
IndoCulture consists of 12 topics and 66 fine-grained subtopics. Below is a list of topics along with their detailed subtopics. The numbers following the topic indicate the total instances required to be written by one worker (with a total of 150 per worker).
  1. 1.
    Food (22): breakfast (2); lunch (3); dinner (2); snacks (2); food souvenir (3); traditional foods and beverages (5); eating habit (1); cutlery (1); cooking ware (1), fruits (2).
  2. 2.
    Wedding (20): traditions before marriage (3); traditions when getting married (3); traditions after marriage (3); men’s wedding clothes (2); women’s wedding clothes (2); invited guests (2); wedding location (1); foods at a wedding (2); gifts brought to weddings (2).
  3. 3.
    Family relationship (13): relationships within the main family (3); relationships in the extended family (3); relations with society/neighbors (5); clan/descendant system (2).
  4. 4.
    Pregnancy and kids (16): traditions during pregnancy (4); traditions after birth (2); how to care for a newborn baby (2); how to care for toddlers (2); how to care for children (2); how to care for teenagers (2); parents and children interactions as adults (2).
  5. 5.
    Death (10): when death occurs (2); the process of caring for a corpse (2); traditions after the body is buried (2); the clothes of the mourners (2); inheritance matters (2).
  6. 6.
    Religious holiday (12): traditions before religious holidays (2); traditions leading up to religious holidays (4); traditions during religious holidays (5); traditions after religious holidays (1).
  7. 7.
    Agriculture (6): what to plant (2); traditions when planting (2); harvest (2).
  8. 8.
    Fisheries and trade (7): traditions of taking care of livestock/fish (5); buying and selling traditions (2)
  9. 9.
    Art (16): musical instruments (3); folk songs (3); traditional dances (3); use of art at certain events (5); poetry or similar literature (2)
  10. 10.
    Traditional games (5): game types; (3), location played (2).
  11. 11.
    Daily activities (10): morning activities (1); afternoon activities (1); evening activities (1); leisure activities (3); house, household and transportation (4).
  12. 12.
    Socio-religious aspects of life (13): regular religious activities (2); mystical things (2); traditional ceremonies (1); lifestyle (3); self care (1); traditional medicine (3); traditional saying (1).
Instance Writing
For each instance, workers craft two culturally relevant sentences that align with the predefined subtopic. The first sentence serves as the premise context, while the last sentence acts as the correct answer. Subsequently, the annotator generates two additional plausible sentences as distractors by modifying cultural objects or activities from the correct sentence. Workers are given a period of two months to complete the task.
Two Stages of Quality Control
In stage 1, we implement quality control by pairing two annotators from the same province. Each annotator is tasked with answering a set of questions prepared by the other annotator, and vice versa. During this phase, the annotator is presented with a premise sentence and three shuffled options. They are allowed to search for the answer from any source if they are unsure. Instances that are incorrectly answered by the second annotator are discarded, as we hypothesize that these instances may contain incorrect answers or possess a level of ambiguity. Additionally, annotators are required to identify whether the instance is province-specific (binary annotation: True/False), meaning it is uniquely relevant in their province and not in others.
In stage 2 of quality control, the first two authors of this paper perform post-editing on the data that has passed the first stage of quality control. We first focus on correcting the linguistic aspects of the text, which includes checking for spelling typos. Although the data is written in Indonesian, some annotators may use dialects or be influenced by the structure or style of regional languages. In these cases, we correct the text to adhere to correct Indonesian grammar.
To maintain the quality of IndoCulture, we rigorously filter instances that display (1) poor writing, as enhancing their quality substantially is difficult; (2) obvious answer options, which allow for easy guessing of the correct choice without understanding the cultural contexts; and (3) ambiguous contexts, where all options are equally valid as the correct answer. For example, in a topic about breakfast, the three options might include one traditional food alongside two other very commonly consumed foods in Indonesia.
Furthermore, we manually verify the province-specific annotation for each instance using the Google search engine. We annotate whether the instance pertains to national-level culture. If the example is specific to a province, it will be annotated as uncommon in national culture, and vice versa.
3.2 Data Statistics
After the annotation process, we initially collected 3,162 instances. [^4] Although we requested each annotator to produce 150 instances, not all were able to complete their allotted tasks within the given timeframe. Unfortunately, we were unable to find additional candidates from the same local province to address the data deficiencies Winata et al. (2023). In stage 1 of quality control, the initial pool of 3,162 instances was reduced to 2,801 instances. Our thorough stage 2 of quality control further reduced the sample to 2,429 final high-quality samples. The data distribution of IndoCulture per province is presented in Table 1. Approximately three-quarters of IndoCulture instances contain province-specific content, with an average length of around 35 words. IndoCulture covers multiple topics, as illustrated in Figure 2.
Figure 2: Topic distribution in IndoCulture.
Province # province $\mu$(word) $\mu$(char)
specific (%)
Aceh 246 70.7 28.0 175.9
North Sumatra 234 83.8 36.8 246.0
West Sumatra 299 74.6 39.6 261.4
West Java 231 58.0 37.5 244.8
Central Java 171 66.7 39.3 260.5
East Java 233 69.5 46.0 310.4
Bali 241 76.3 33.3 216.1
NTT 103 72.8 31.8 203.6
South Borneo 233 83.7 33.3 226.0
South Sulawesi 185 90.3 33.6 227.8
Papua 253 88.1 37.3 245.0
All 2429 76.0 NA NA
Table 1: Overall statistics of IndoCulture by province.
4 Experiments
4.1 Set-Up
We focus on evaluating 23 language models in zero-shot settings: (1) 15 open-source multilingual language models including BLOOMZ Muennighoff et al. (2022), mT0 Muennighoff et al. (2022), LlaMA2 Touvron et al. (2023), Bactrian-X Li et al. (2023a) in different sizes, (2) 2 South East Asian language models including SeaLLM Nguyen et al. (2023b), and SeaLion Singapore (2023), (3) 4 Indonesian-centric language models, including IndoBART Cahyawijaya et al. (2021), IndoGPT Cahyawijaya et al. (2021), Merak Ichsan (2023), Komodo Owen et al. (2024), (3) 2 closed-source models, including ChatGPT: gpt-3.5-turbo Ouyang et al. (2022) and GPT–4: gpt-4-0613 OpenAI (2023). Please refer to Appendix A for further details.
Initially, we evaluate the effectiveness of sentence completion and multiple-choice question strategies in predicting the correct options. In both scenarios, we conduct benchmarks across three distinct location contexts. Formally, given a premise $s$, three candidate options $c_{1}$, $c_{2}$, $c_{3}$, and location $l\in\{\texttt{none},\texttt{Indonesia},p_{j}\}$, for sentence completion, we select the correct option based on:
[…]
Here, $\text{concat}(s,c)$ denotes the concatenation of premise $s$ and candidate option $c$, separated by a space. In the case of multiple-choice questions, we devise a template for the prompt question and determine the answer by selecting the option with the highest probability among letters A, B, and C. Figure 3 displays the prompt templates in Indonesian and English for both sentence completion and multiple-choice question.
Figure 3: Templates for sentence completion and multiple-choice questions prompts.
4.2 Results
Model (#parameter) Completion MCQ
$l=\text{None}$ $l=\text{Ind}$ $l=\text{Prov}$ $l=\text{None}$ $l=\text{Ind}$ $l=\text{Prov}$
Human 100.0 100.0
Random 33.3 33.3 33.3 33.3 33.3 33.3
BLOOMZ (560M) 37.5 35.3 35.3 32.1 32.0 32.1
BLOOMZ (1.1B) 36.2 36.9 37.2 32.0 32.0 32.0
BLOOMZ (3B) 38.5 40.5 41.2 45.9 47.6 48.6
BLOOMZ (7B) 40.9 43.8 44.4 48.8 49.8 49.7
[…]
LLaMA2 (7B) 36.7 37.0 37.2 40.0 39.0 37.7
LLaMA2-chat (7B) 36.9 36.9 37.4 40.5 41.1 40.3
LLaMA2 (13B) 39.4 39.7 39.9 46.9 46.9 46.6
LLaMA2-chat (13B) 38.2 38.4 38.9 47.2 49.2 49.2
[…]
IndoBART (132M) 41.9 41.1 41.7 32.1 32.0 32.2
IndoGPT (117M) 42.6 42.1 42.7 33.7 34.1 35.0
Merak (7B) 40.2 40.6 42.6 51.5 52.5 52.5
SeaLLM (7B) 39.1 39.7 41.1 51.7 52.4 52.5
SEA-LION (7B) 38.8 38.7 39.6 33.4 32.8 33.0
Komodo (7B) 45.1 45.4 46.1 37.6 35.1 36.1
GPT–3.5 (NA) 58.4 60.2 61.7
GPT–4 (NA) 68.8 71.5 75.8
Table 2: Zero-shot accuracy across various models and settings. “MCQ” refers to the multiple-choice question method, and $l$ denotes the location as additional context. The bold numbers highlight the highest score within each model group.
Overall observation
[…]
Multiple-choice question method is generally better.
Our findings suggest that the multiple-choice question method tends to outperform the sentence completion method, with exceptions noted for BLOOMZ (560M, 1.1B), IndoBART, IndoGPT, and Komodo. Interestingly, within the sentence completion method, the small Indonesian-centric language model, IndoGPT, surpasses the performance of almost all large multilingual models and remains competitive with BLOOMZ (7B). However, IndoGPT’s performance experiences a significant decline in the multiple-choice question method, with a notable margin of nearly 8 to 10 points. While Komodo (7B) achieves the highest accuracy with the sentence completion method, its performance also decreases by 10-12 points in the multiple-choice method. This discrepancy could potentially be attributed to differences in the nature of language model training and instruction-tuning.
Impact of location context on model performance
Our investigation reveals that incorporating various levels of location granularity has a noticeable effect on zero-shot performances, especially models with larger parameter sizes. Detailed location contexts notably enhance the accuracy of BLOOMZ (7B), LlaMA2 (13B), Merak (7B), SeaLLM (7B), Komodo (7B), GPT–3.5, and GPT–4. For instance, in GPT–4, the accuracy gap between $l=\texttt{none}$ and $l=\texttt{Indonesia}$ is 2.7, and this gap further increases by 7 points when $l=\texttt{Prov}$ is assigned.
4.3 Analysis
Given the superior performance of models with large parameter sizes using the multiple-choice question method and location $l=\texttt{Prov}$, we employ these settings for analysis.
Province Merak GPT–3.5 GPT–4
$\neg$¬PS PS $\neg$¬PS PS $\neg$¬PS PS
Aceh 59.7 53.4 79.2 65.5 93.1 73.6
North Sumatra 58.3 46.9 63.9 57.2 75.0 73.2
West Sumatra 50.0 41.3 63.2 48.4 85.5 63.7
West Java 65.3 58.8 84.2 71.8 92.6 81.7
Cental Java 58.8 47.1 67.6 52.9 82.4 72.9
East Java 57.4 37.6 72.2 46.5 87.0 63.4
Bali 78.9 65.6 87.7 81.7 93.0 89.4
NTT 64.3 52.0 57.1 49.3 85.7 68.0
South Borneo 65.7 46.4 68.6 53.6 77.1 67.4
South Sulawesi 53.3 47.9 66.7 55.8 80.0 70.6
Papua 76.7 55.2 76.7 62.8 90.0 71.3
Table 3: Top-3 model accuracy by province. "PS" indicates instances containing province-specific contexts, while "$\neg$¬PS" indicates otherwise. The green and red cells indicate the top three and bottom three scores, respectively.
Results by province
Table 3 highlights that the top 3 performing LLMs exhibit a nuanced understanding of cultures within Indonesian provinces, particularly excelling in the cultures of West Java and Bali compared to other provinces. GPT–4, for instance, achieves the best accuracy of 93%, while Merak and GPT–3.5 also demonstrate strong performance in Bali and West Java, with accuracies ranging between 55% and 88%. In other provinces like West Sumatra and South Borneo, the models typically exhibit poorer performance. Specifically, for GPT-3.5, the performance gap compared to Bali ranges from 19 to 33%. This highlights the presence of cultural biases and a lack of inclusivity in model reasoning abilities, likely stemming from the distribution of training data. The proximity of West Java to Jakarta (Indonesia’s capital) and Bali’s global status as a tourism destination may contribute to the abundance of textual data on these two cultures.
We also note a consistent disparity between non-province and province-specific contexts across all models, with models generally finding non-province contexts easier to comprehend. On average, this gap ranges from 12 to 13 points for the three models, highlighting the challenge posed by province-specific content and emphasizing the significant influence of location context on the reasoning ability of LLMs.
Results by topic
Table 4 shows the accuracy of the top 3 performing models across topics. Similar to Table 3, the models perform better in province-specific context for all topics, with death customs being an exception. For province-specific contexts, GPT–4 excels in themes of wedding, death, and arts while for non-specific contexts, GPT–4 achieves accuracy more than 90.0 for family relationship, religious holidays, and farm and fishery.
Province Merak GPT–3.5 GPT–4
$\neg$¬PS PS $\neg$¬PS PS $\neg$¬PS PS
Food 56.9 51.0 65.9 62.1 79.3 63.5
Wedding 66.7 64.3 77.6 56.0 92.8 78.6
Family relationship 65.5 54.5 72.5 53.9 93.1 72.7
Pregnancy and kids 67.9 42.9 69.2 50.6 82.1 72.7
Death 38.5 55.2 60.9 62.5 76.9 82.5
Religious holidays 68.5 48.8 79.3 68.8 94.4 64.6
Agriculture 70.5 52.3 71.4 62.1 88.5 76.1
Farm and fishery 77.8 51.0 77.8 51.9 94.4 75.8
Art 58.7 50.2 85.2 61.1 78.3 77.3
Traditional game 60.0 50.2 80.3 59.8 88.0 66.9
Daily activities 55.1 39.6 66.0 64.8 87.8 63.0
Socio-religious 57.1 44.3 79.6 61.1 85.7 64.6
Table 4: Top-3 model accuracy by topic. "PS" indicates instances containing province-specific contexts, while "$\neg$¬PS" indicates otherwise. The green and red cells indicate the top three and bottom three scores, respectively.
Does language affect model performance?
We automatically translated IndoCulture to English using the Google Translate API [^5] Accessed in March 2024. and used the English prompt in Figure 3 to evaluate the models. All model performances in English text dropped except for LLaMA-2 and Merak. This could be attributed to two reasons. First, LLaMA-2 is an English-centric model, and Merak is pretrained from LLaMA-2. Second, the performance drop in other models could be caused by translation errors. We further investigated this with 100 random samples and found that 81 samples had acceptable translations. We observed translation errors such as pronoun mismatches, inaccurate proverb translations, and inaccurate translations of local terms, such as “pupuik” translated as “fertilizer”.666pupuik is a traditional musical instrument in West Sumatra. It is worth noting that the word pupuik closely resembles the word pupuk in Indonesian, which means fertilizer. To better understand the cultural gap in language models, we followed the approach of Liu et al. (2023a) to manually correct the translations and reevaluate the models. We found that GPT-4’s performance over the 100 random samples was 77.0 for the original Indonesian text, 68.0 for the English machine-translated text, and 72.0 for the English translation fixed by humans.
[…]
Results by fine-grained culture elements
We tasked two expert workers with annotating 200 random samples from IndoCulture based on six cultural elements, as derived from Axtell and Fornwald (1998); Williams (2014). [^7] https://www.languageeducatorsassemble.com/elements-of-culture [^8] https://pressbooks.howardcc.edu/soci101/chapter/3-2-the-elements-of-culture/ While these elements may not encompass every cultural aspect, we contend that they cover the most prominent or pivotal elements, including: (1) symbols (material or non-material objects representing meaning), (2) artifacts (material or non-material objects produced by society), (3) values and beliefs (principles, ideas, and concepts assumed to be ideal and correct in society), (4) norms (rules guiding values and beliefs), (5) language, (6) rituals (established procedures and ceremonies), and (7) other, for examples that do not fit into any of the defined elements. This annotation is a multi-label task, and the average Kappa score across the cultural elements is 0.56, with each ranging from 0.4 to 0.75. These scores indicate moderate to substantial agreement.
Table 5 displays the distribution of each cultural element in our dataset, along with the performance breakdown across Merak, GPT-3.5, and GPT-4. Among the 200 random samples, we observe that 42.5% of our data contains artifacts, 37.5% norms, and 30% rituals. Only 4% of the data pertains to symbols, while 7.5% belongs to the other category. Merak exhibits lower performance in norms, with a 24% decrease compared to values and beliefs. Conversely, GPT-3.5 and GPT-4 perform better in norms, with the performance gap in values and beliefs being only $-4$ points. Additionally, language poses a challenge in Merak, with only 38% accuracy, while GPT-3.5 and GPT-4 achieve 60% and 72% accuracy, respectively.
Cultural element (%) Merak GPT–3.5 GPT-4
Symbols (4) 50.0 58.9 70.8
Artefacts (42.5) 55.3 59.1 67.8
Values and Beliefs (10.5) 61.9 57.5 69.8
Norms (37.5) 38.7 63.2 73.6
Language (19.5) 38.5 60.2 72.0
Ritual (30) 53.3 60.0 70.7
Other (7.5) 66.7 58.9 69.2
Table 5: Accuracy comparison of Merak, GPT–3.5, and GPT–4 across 200 random samples, categorized by cultural elements. The numerical value following each cultural element indicates its proportion within the samples.
Can the model provide a reasonable explanation to support the answer?
Figure 5: Performance comparison between Merak (7B) and GPT–4 based on text generation output. “Answer (T)” indicates that the generated answer is true, while ”Exp(F)” denotes that the answer explanation is false.
We conduct a manual investigation of the text generation output for Merak (7B) and GPT–4 across 200 random samples. This involves manually examining the generated answer along with its explanation. To obtain explanation, we modify the Indonesian prompt in Figure 3 by adding the string Jelaskan jawabanmu! ("Explain your answer!"). Our annotation process is binary, categorizing explanations as either True or False. We label an explanation as False if it is absent, contains hallucinations, or provides inaccurate information. [^9] We use Google search engine to verify the correctness of the explanation.
As anticipated, there is a significant drop in accuracy for Merak (7B) from 52.5% (as shown in Table 2) to 29.5%. This discrepancy underscores the limitations of relying solely on token probability to assess the true capability of a language model. Interestingly, only 4.5% of the samples are answered correctly with the appropriate explanation by Merak, despite it being the top performer among the open-source models. GPT–4 achieves an accuracy of 69.5%, which is 5 points lower than indicated in Table 2, possibly due to a smaller sample size and biases in regular expressions. However, GPT–4 provides better explanations, with 51.5% of the samples correctly answered along with the appropriate explanation.
5 Discussion
The recent study Wang et al. (2024) has demonstrated that evaluating language models using token probability in multiple-choice question types does not align well with the generated text. This discrepancy is reported to be more pronounced in models fine-tuned on conversational or safety data. In response to this issue, we have conducted a manual evaluation on 200 random samples (in Section 4.3) and found that the performance in generated texts, especially for Merak, the best open-source language model, is deteriorating. However, this issue is less apparent in GPT-4, as the initial evaluation is based on regular expressions. Conducting manual evaluation on all data and models is expensive, and we plan to address this issue in future works. This work primarily focuses on introducing a novel dataset meticulously constructed for evaluating cultural commonsense reasoning within the Indonesian context, including preliminary evaluation results based on the standard methods used in previous studies OpenAI (2023); Touvron et al. (2023); Koto et al. (2023); Li et al. (2023b); Koto et al. (2024).
6 Conclusion
IndoCulture is a cultural commonsense reasoning dataset encompassing the diversity of Indonesian cultures, spanning from Aceh province in the west to Papua province in the east. Through collaboration with local individuals across eleven provinces and rigorous quality control measures, we introduce IndoCulture for the purpose of evaluating language models. Our findings reveal that large language models, whether Indonesian-centric or multilingual, demonstrate a limited understanding of Indonesian culture. Notably, incorporating location as additional context significantly enhances model performance, particularly in GPT–3.5 and GPT–4.
Limitations
IndoCulture is specifically designed to explore the influence of geographical location on cultural commonsense reasoning, with a focus on the present time. It does not consider temporal aspects. Our dataset was created in the year 2023, recognizing that cultures may evolve over time, as discussed by Mesoudi (2016).
Furthermore, as demonstrated in Section 4.3, a significant portion of IndoCulture comprises cultural elements such as artifacts, norms, and rituals. Symbols, values and beliefs, and language represent smaller proportions, ranging from 4% to 20%. We encourage future research to further explore these cultural elements and to expand the geographical coverage beyond the eleven provinces studied in this paper.
Ethical Considerations
IndoCulture is licensed under a Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License. [^10] https://creativecommons.org/licenses/by-nc-sa/4.0/ Our data is intended for academic research and non-commercial purposes. Workers are compensated above the minimum monthly salary in Indonesia and are fully aware that the data will be released to the public. It is important to note that no private or sensitive information of the workers is included in IndoCulture.
Appendix A Model Artifacts
Models (#parameters) Source
BLOOMZ (560M) bigscience/bloomz-560m
BLOOMZ (1.1B) bigscience/bloomz-1b1
BLOOMZ (1.7B) bigscience/bloomz-1b7
BLOOMZ (3B) bigscience/bloomz-3b
BLOOMZ (7.1B) bigscience/bloomz-7b1
[…]
LLamA2-chat (7B) meta-llama/Llama-2-7b-chat
LLamA2 (13B) meta-llama/Llama-2-13b
LLamA2-chat (13B) meta-llama/Llama-2-13b-chat
[…]
IndoGPT (117M) indobenchmark/indogpt
Merak (7B) Ichsan2895/Merak-7B-v5-PROTOTYPE1
SeaLLM (7B) SeaLLMs/SeaLLM-7B-v2
SEA-LION (7B) aisingapore/sea-lion-7b
Komodo (7B) Yellow-AI-NLP/komodo-7b-base
Table 6: With the exception of GPT-3.5 and GPT-4, all the models used in this study were sourced from Huggingface Wolf et al. (2020).

### v2

@@
 Abstract
-Although commonsense reasoning is greatly shaped by cultural and geographical factors, previous studies on language models have predominantly centered on English cultures, potentially resulting in an Anglocentric bias.
+Although commonsense reasoning is greatly shaped by cultural and geographical factors, previous studies have predominantly centered on cultures grounded in the English language, potentially resulting in an Anglocentric bias.
 In this paper, we introduce IndoCulture, aimed at understanding the influence of geographical factors on language model reasoning ability, with a specific emphasis on the diverse cultures found within eleven Indonesian provinces.
-In contrast to prior works that relied on template Yin et al.
+In contrast to prior work that has relied on templates Yin et al.
 (2022) and online scrapping Fung et al.
-(2024), we create IndoCulture by asking local people to manually develop the context and plausible options based on predefined topics.
-Evaluations of 23 language models reveal several insights: (1) even the best open-source model struggles with an accuracy of 53.2%, (2) models often provide more accurate predictions for specific provinces, such as Bali and West Java, and (3) the inclusion of location contexts enhances performance, especially in larger models like GPT-4, emphasizing the significance of geographical context in commonsense reasoning.
-[^1] IndoCulture can be accessed at https://huggingface.co/datasets/indolem/IndoCulture
+(2024), we create IndoCulture by asking local people to manually develop a cultural context and plausible options, across a set of predefined topics.
+Evaluation of 27 language models reveals several insights: (1) the open-weight Llama–3 is competitive with GPT–4, while other open-weight models struggle, with accuracies below 50%; (2) there is a general pattern of models generally performing better for some provinces, such as Bali and West Java, and less well for others; and (3) the inclusion of location context enhances performance, especially for larger models like GPT–4, emphasizing the significance of geographical context in commonsense reasoning.
+[^1] IndoCulture is available from https://huggingface.co/datasets/indolem/IndoCulture
 1 Introduction
@@
 The highlighted regions in the map represent the provinces examined in IndoCulture.
-We offer examples from Aceh, South Borneo, and Papua, with correct answers indicated in bold.
-English translations are for illustrative purposes.
-Reasoning abilities of multilingual language models are frequently evaluated using English texts, potentially amplifying an Anglocentric bias toward English culture and leading to less inclusive models Thomas (1983); Ponti et al.
+We present examples from Aceh, North Sumatra, and Papua, with three plausible options and correct answers indicated in bold.
+English translations are provided for illustrative purposes.
+The reasoning abilities of multilingual language models are frequently evaluated using English texts, potentially amplifying an Anglocentric bias toward culture grounded in the English language, and leading to less inclusive models Thomas (1983); Ponti et al.
 (2020).
-Cultures, however, vary significantly from one location to another and profoundly shape the way a speaker of a language reasons Hershcovich et al.
+Cultures, however, vary significantly from one location to another and profoundly shape the way speakers of a language reason Hershcovich et al.
 (2022).
-Recent evaluations of models reasoning ability OpenAI (2023); Touvron et al.
-(2023); Jiang et al.
-(2023); Sengupta et al.
+Recent evaluations of models’ commonsense reasoning ability OpenAI (2023); Sengupta et al.
 (2023); Liu et al.
-(2023b) conducted on the English datasets, such as PIQA Bisk et al.
-(2020a), HellaSwag Zellers et al.
-(2019), Social IQA Sap et al.
-(2019), often overlook geographical aspects, thereby risking cultural bias.
-Culture is a multifaceted concept encompassing the way of life Giddens and Sutton (2021), including our thoughts and actions Macionis John .
-It includes tangible elements like food art, and clothing, as well as intangible aspects such as ideas, values, attitudes, and norms.
+(2023) have been conducted on English datasets such as Social IQA Sap et al.
+(2019) and PIQA Bisk et al.
+(2020), and thus often overlook geographical aspects, thereby risking cultural bias.
+Culture is a multifaceted concept encompassing the way of life Giddens and Sutton (2021), including our thoughts and actions Macionis (2012).
+It includes tangible elements like food, art, and clothing, as well as intangible aspects such as ideas, values, attitudes, and norms.
 Culture is shaped by geographical location and ethnicity, influencing the commonsense reasoning of people within a region.
-For example, in Indonesia, it is culturally acceptable to eat rice with your hands but it is considered unusual to use a fork.
+For example, in Indonesia, it is culturally acceptable to eat rice with your hands but it is considered unusual to use chopsticks.
 Similarly, at traditional Indonesian weddings, it is common to sit on the floor while eating, whereas this practice is less common in Australia.
@@
 (2024); Putri et al.
-(2024) overlooked the geographical nuance and the rich diversity of Indonesian cultures.
+(2024) have explicitly addressed the geographical nuances and rich diversity of Indonesian cultures.
 This paper introduces IndoCulture, a novel dataset to evaluate cultural reasoning in eleven Indonesian provinces, manually developed by local people in each province based on predefined topics.
-In prior works, cultural reasoning has primarily relied on datasets constructed through templates Yin et al.
+In prior work, cultural reasoning has primarily relied on datasets constructed through templates Yin et al.
 (2022), and online scraping Nguyen et al.
-(2023a); Fung et al.
+(2023); Fung et al.
 (2024).
-While these studies offer valuable insights, they may be susceptible to training data contamination when utilized to assess large language models (LLMs).
+While these studies offer valuable insights, they may be susceptible to training data contamination when used to assess large language models (LLMs).
 For instance, Fung et al.
-(2024) reported a zero-shot accuracy of 92% when using ChatGPT to evaluate low-resource data.
-Figure 1 illustrates the eleven provinces examined in IndoCulture including Aceh, North Sumatra, West Sumatra, West Java, Central Java, East Java, Bali, South Borneo, East Nusa Tenggara (NTT), South Sulawesi, and Papua.
+(2024) reported a zero-shot accuracy of 92% when using ChatGPT Ouyang et al.
+(2022) to evaluate low-resource data.
+IndoCulture contains cultural commonsense knowledge data from eleven provinces in Indonesia (blue colored in Figure 1), namely Aceh, North Sumatra, West Sumatra, West Java, Central Java, East Java, Bali, South Borneo, East Nusa Tenggara (NTT), South Sulawesi, and Papua.
+These provinces span breadth of Indonesia, each representing a major island in the country, with the addition of Bali and NTT.
+Figure 1 also shows three examples in IndoCulture for three provinces: Aceh, North Sumatra, and Papua.
 [^2] Although Papua consists of six provinces, for the purpose of this study, we treat it as a single entity (referred to as Papua) due to the relatively recent establishment of most of these provinces.
-These provinces span from the eastern to the western parts of Indonesia, each representing a major island in the country, with the addition of Bali and NTT.
-Figure 1 also shows three examples in IndoCulture for three provinces: Aceh, South Borneo, and Papua.
-The first two examples focus on cultural artifacts: the traditional wedding dress from Aceh and the Bagandut dance from South Borneo.
-The second example also examines cultural symbols.
-The third example focuses on cultural beliefs and norms regarding pregnancy in Papua.
+The first example focuses on cultural artifact, specifically the traditional wedding dress from Aceh.
+The second example examines family relationships while the third example focuses on cultural beliefs and norms regarding pregnancy in Papua.
+Dataset Size Data Construction Method Cultural? Location? #province #topic
+IndoCulture (ours) 2,429 Manually built and validated by native 11 66
+COPAL-ID Wibowo et al. (2024) 559 Manually built and validated by native
+MAPS Liu et al. (2024) 371 LLM generation & human generation 1
+ID-CSQA Putri et al. (2024)* 4,416 LLM generation & human generation 5
+BLEnD Myung et al. (2024) 1,000 Template, translation, human validation 1 6
+IndoCloze Koto et al. (2022) 2,335 Manually built and validated by native
+XCOPA Ponti et al. (2020) 600 Translated from English data
+XStoryCloze Lin et al. (2022) 1,872 Translated from English data
+Table 1: Comparison of IndoCulture with other cultural knowledge and reasoning datasets containing instances in Indonesian.
+The metadata includes Size (number of Indonesian instances), Cultural?
+(whether the data considers cultural nuances), Location?
+(whether the data includes fine-grained location information, such as provinces, as context), #province (number of Indonesian provinces covered), and #topic (number of fine-grained topics covered).
+* indicates the dataset involves question generation with less emphasis on reasoning.
 Can large language models effectively reason based on the diverse cultures of Indonesia?
-To capture the rich diversity of Indonesian cultures, we meticulously predefined 12 fine-grained topics as guidelines for data construction.
+To capture the rich diversity of Indonesian cultures, we predefined 12 fine-grained topics as guidelines for data construction.
 Figure 2 displays the topic distribution in IndoCulture, with the majority focusing on food, weddings, art, pregnancy and children, and family relationships.
-Additionally, we also pose the question: Is there an influence of geographical location on commonsense reasoning by language models?
+Additionally, we also pose the question: Is there any influence of geographical location on the commonsense reasoning of language models?
 We address these questions through comprehensive experiments across different language models, incorporating several levels of location granularity as additional context in the prompt.
@@
   • We present IndoCulture, a high-quality cultural reasoning dataset in the Indonesian language, covering eleven provinces of Indonesia and twelve fine-grained cultural topics.
-Our dataset has 2,429 instances, and was meticulously developed by local people with rigorous quality controls in place.
-  • We assess 15 open-source multilingual models, 6 open-source Indonesian-centric models, and 2 closed-source models.
-Although local individuals can answer all questions correctly (i.e., 100% accuracy), we observed that all open-source models struggle to comprehend Indonesian cultures, with the highest performance reaching only 53.2%.
-  • We conduct a thorough analysis over various dimensions: (1) models performance for each province and topics, (2) the influence of different granularities of location contexts (i.e., none, province, country), (3) model performance over English translation, (4) analysis over model explanation behind its answer.
+Our dataset has 2,429 instances, and was developed by local people with rigorous quality controls in place.
+  • We assess 19 open-weight multilingual models, 6 open-weight Indonesian-centric models, and 2 closed-weight models.
+Although local individuals can answer all questions correctly (i.e., 100% accuracy), most open-weight models struggle to comprehend Indonesian cultures.
+Interestingly, we observed that Llama–3 Dubey et al.
+(2024) is competitive with GPT–4 OpenAI (2023).
+  • We conduct a thorough analysis over various dimensions: (1) model performance for each province and topic; (2) the influence of different granularities of location context (i.e., none, province, country); (3) model performance over English translations; and (4) analysis of model explanations for a given answer.
 2 Related Work
 Commonsense Reasoning in English
-Many studies have focused on commonsense reasoning in English, often overlooking considerations of cultures and geographical locations.
-Early works include the Winograd Schema Challenge Levesque et al.
+Many studies have focused on commonsense reasoning in English, often overlooking considerations of culture and geographical location.
+Early work included the Winograd Schema Challenge Levesque et al.
 (2012) and WinoGrande Sakaguchi et al.
-(2021) for pronoun coreference resolution tasks.
+(2021) for pronoun coreference resolution.
 Other research areas include reasoning based on cause-effect relationships Roemmele et al.
 (2011), physical activities Bisk et al.
-(2020b), social interaction Sap et al.
+(2020), social interactions Sap et al.
 (2019), cloze story completion Mostafazadeh et al.
 (2016), sentence completion Zellers et al.
-(2019), numerical Lin et al.
-(2020), and temporal Qin et al.
+(2019), numerical reasoning Lin et al.
+(2020), and temporal reasoning Qin et al.
 (2021).
-Additionally, pretrained language models are employed in other works to extract structured commonsense knowledge by providing seed words Davison et al.
-(2019), and using code language model Madaan et al.
+Additionally, pretrained language models have been employed in other work to extract structured commonsense knowledge by providing seed words Davison et al.
+(2019), and using code language models Madaan et al.
 (2022).
 Cultural Commonsense Reasoning with Geographical Contexts
-Previous studies have explored commonsense reasoning with geographical contexts.
-Shwartz (2022) investigated time perception (e.g., morning and night hours) across different locations, while Yin et al.
-(2022) delved into cultural knowledge of language models across five countries by masking and template strategies.
-Other works have focused on automatically extracting cultural knowledge from various sources, including Wikipedia Fung et al.
+Previous studies have explored commonsense reasoning with geographical context.
+Shwartz (2022) investigated time perception (e.g., morning and night) across different locations, while Yin et al.
+(2022) examined cultural knowledge of language models across five countries using datasets built from templates and translations.
+Other work has focused on automatically extracting cultural knowledge from various sources, including Wikipedia Fung et al.
 (2024), conversations Fung et al.
 (2023), and Common Crawl Nguyen et al.
-(2023a), incorporating location contexts with the assistance of large language models.
+(2023), incorporating location context with the assistance of large language models.
 Related, Ziems et al.
-(2023) created a knowledge bank for situational norms, utilizing English-speaking Mechanical Turk annotators and incorporating country taxonomy.
-Unlike these works, IndoCulture specifically concentrates on cultural reasoning within Indonesian provinces, developed and validated manually by local people (experts).
-Compared to the automatic method and English-speaking crowdsource workers for data construction, IndoCulture arguably contains less noise, and is free from training data contamination of large language models (LLMs).
+(2023) created a knowledge bank for situational norms, using English-speaking Mechanical Turk annotators and incorporating a country taxonomy.
+Unlike this work, IndoCulture specifically concentrates on cultural reasoning across Indonesian provinces, developed and validated manually by local people (experts).
+Compared to the automatic method and English-speaking crowd workers for data construction, IndoCulture arguably contains less noise, and is free from the training data contamination of large language models (LLMs).
 Commonsense Reasoning with Indonesian contexts
+Table 1 shows a comparison of IndoCulture with other Indonesian datasets for cultural knowledge and reasoning evaluation.
 Commonsense reasoning in Indonesian language models has been studied using translated English–Indonesian datasets, such as XCOPA Ponti et al.
@@
 (2022).
-However, these datasets potentially introduce a cultural bias toward English culture.
+However, these datasets potentially introduce a cultural bias toward culture grounded in the English language.
 IndoCloze Koto et al.
-(2022) is the first commonsense reasoning dataset in Indonesian, developed by native Indonesian workers following cloze story completion framework Mostafazadeh et al.
+(2022) was the first commonsense reasoning dataset in Indonesian, developed by native Indonesian workers following the cloze story completion framework Mostafazadeh et al.
 (2016).
-IndoMMLU Koto et al.
-(2023), a benchmark for assessing knowledge in the Indonesian context, includes questions pertaining to Indonesian culture.
-However, its primary focus lies on testing proficiency in the local language rather than reasoning abilities.
-In a contemporary study, Wibowo et al.
+However, IndoCloze lacks local cultural nuances and fine-grained geographical context.
+Wibowo et al.
 (2024) followed the COPA framework Roemmele et al.
@@
 In other work, Putri et al.
-(2024) studied the capability of LLMs in generating questions with cultural norms, for both general Indonesian and specific Sundanese contexts.
-In comparison, IndoCulture offers broader coverage across eleven provinces, thereby fostering greater inclusivity for local communities in Indonesia.
+(2024) studied the capability of LLMs in generating questions with cultural norms, for both general Indonesian and specific Sundanese contexts, while Liu et al.
+(2024) used proverbs and LLMs to generate conversational data.
+In contemporary work, Myung et al.
+(2024) released BLEnD, a large-scale cultural knowledge dataset, built using templates, translation, and human validations, covering the West Java province in Indonesia.
+BLEnD specifically focuses on short-answer questions, limiting its capacity for reasoning evaluation.
+Unlike most other datasets that do not consider geographical factors, IndoCulture has broad coverage across eleven provinces, thereby providing greater inclusivity for local communities in Indonesia.
 3 IndoCulture
-As illustrated in Figure 1, IndoCulture is a sentence completion task in Indonesian language featuring a one-sentence premise, three plausible options, and one correct option to evaluate reasoning ability and cultural knowledge across eleven Indonesian provinces.
-In total, there are 2,429 instances, with the distribution across provinces detailed in Table 1.
+As illustrated in Figure 1, IndoCulture is a sentence completion task in the Indonesian language featuring a one-sentence premise, three plausible options, and one correct option to evaluate reasoning ability and cultural knowledge across eleven Indonesian provinces.
+While sentence completion tasks are straightforward for humans, answering IndoCulture requires machines to engage in cultural reasoning to logically conclude which of the three options is logically consistent with the first sentence Huang and Chang (2023).
+The dataset includes a total of 2,429 instances.
 3.1 Data Construction
-IndoCulture is constructed manually by human and verified through a two-step process.
+IndoCulture was constructed manually by humans, and verified through a two-step process.
 Worker Recruitment
-Culture often arises from the shared experiences, traditions, and beliefs of a specific group over time, often closely intertwined with native populations.
-Keeping this in mind, we engaged individuals from various provinces across Indonesia to assist in preparing data for the IndoCulture benchmark.
-During the recruitment, we presented a few examples of the intended IndoCulture data and requested each candidate to generate similar instances tailored to the context of their respective provinces.
+Culture generally arises from the shared experiences, traditions, and beliefs of a specific group over time, often closely intertwined with native populations.
+With this in mind, we engaged individuals from various provinces across Indonesia to assist in preparing data for the IndoCulture benchmark.
+During recruitment, we presented a few examples of the intended IndoCulture data and requested each candidate to generate similar instances tailored to the context of their respective provinces.
 From a pool of 58 applicants, we carefully selected 22 expert workers representing 11 provinces (with 2 workers selected per province).
@@
 The age range of our workforce spans from 21 to 35 years old, with educational backgrounds distributed as follows: 3 high school graduates, 14 bachelor’s degree holders, 4 Master’s degree holders, and 1 PhD holder.
-During data construction, each expert worker will fulfill the dual roles of instance writer and quality controller.
-[^3] Each worker is compensated above the monthly minimum wage in Indonesia.
+During data construction, each expert worker fulfilled the dual roles of instance writer and quality controller. Each worker was compensated above the monthly minimum wage in Indonesia.
 Province Selection
-The provinces covered in this study represent diversity of Indonesian cultures.
-Those 11 provinces (in Figure 1) spread across 6 islands and large archipelagos in Indonesia which are inhabited by various tribes who speak different regional languages and adhere to different religions.
+The provinces covered in this study represent the diversity of Indonesian cultures.
+The 11 provinces (in Figure 1) are spread across 6 islands of the Indonesian archipelago, which are inhabited by different ethnic groups who speak different regional languages and adhere to different religions.
 Topic Taxonomy
-IndoCulture consists of 12 topics and 66 fine-grained subtopics.
-Below is a list of topics along with their detailed subtopics.
-The numbers following the topic indicate the total instances required to be written by one worker (with a total of 150 per worker).
+IndoCulture consists of 12 topics and 66 fine-grained subtopics, carefully constructed based on discussions and brainstorming with Indonesian natives.
+The selection of these topics and subtopics was guided based on several criteria and motivations: (1) relevance to Indonesian culture; (2) diversity and coverage; (3) regional representation (e.g., religious holidays); (4) practicality; and (5) expert consultation (i.e., native speaker feedback).
+Compared to the other Indonesian datasets in Table 1, IndoCulture includes a richer array of fine-grained topics.
+Below is a list of the topics along with their detailed subtopics.
+The numbers following each topic indicate the total number of instances required to be written by one worker (with a total of 150 per worker).
   1. 1.
-    Food (22): breakfast (2); lunch (3); dinner (2); snacks (2); food souvenir (3); traditional foods and beverages (5); eating habit (1); cutlery (1); cooking ware (1), fruits (2).
+    Food (22): breakfast (2); lunch (3); dinner (2); snacks (2); food souvenirs (3); traditional foods and beverages (5); eating habits (1); cutlery (1); cooking ware (1), fruit (2).
   2. 2.
@@
   5. 5.
-    Death (10): when death occurs (2); the process of caring for a corpse (2); traditions after the body is buried (2); the clothes of the mourners (2); inheritance matters (2).
+    Death (10): when death occurs (2); the process of dealing with a corpse (2); traditions after the body is buried (2); the clothes of the mourners (2); inheritance matters (2).
   6. 6.
@@
   11. 11.
-    Daily activities (10): morning activities (1); afternoon activities (1); evening activities (1); leisure activities (3); house, household and transportation (4).
+    Daily activities (10): morning activities (1); afternoon activities (1); evening activities (1); leisure activities (3); house, household, and transportation (4).
   12. 12.
-    Socio-religious aspects of life (13): regular religious activities (2); mystical things (2); traditional ceremonies (1); lifestyle (3); self care (1); traditional medicine (3); traditional saying (1).
+    Socio-religious aspects of life (13): regular religious activities (2); mystical things (2); traditional ceremonies (1); lifestyle (3); self care (1); traditional medicine (3); traditional sayings (1).
 Instance Writing
-For each instance, workers craft two culturally relevant sentences that align with the predefined subtopic.
+For each instance, workers were asked to craft two culturally-relevant sentences that align with the predefined subtopic.
 The first sentence serves as the premise context, while the last sentence acts as the correct answer.
 Subsequently, the annotator generates two additional plausible sentences as distractors by modifying cultural objects or activities from the correct sentence.
-Workers are given a period of two months to complete the task.
+These distractors are designed to reflect local cultural contexts, ensuring they are challenging yet unambiguous, and could potentially serve as correct answers in other regional contexts.
+Workers were given a period of two months to complete the task.
 Two Stages of Quality Control
-In stage 1, we implement quality control by pairing two annotators from the same province.
-Each annotator is tasked with answering a set of questions prepared by the other annotator, and vice versa.
-During this phase, the annotator is presented with a premise sentence and three shuffled options.
-They are allowed to search for the answer from any source if they are unsure.
-Instances that are incorrectly answered by the second annotator are discarded, as we hypothesize that these instances may contain incorrect answers or possess a level of ambiguity.
-Additionally, annotators are required to identify whether the instance is province-specific (binary annotation: True/False), meaning it is uniquely relevant in their province and not in others.
-In stage 2 of quality control, the first two authors of this paper perform post-editing on the data that has passed the first stage of quality control.
-We first focus on correcting the linguistic aspects of the text, which includes checking for spelling typos.
-Although the data is written in Indonesian, some annotators may use dialects or be influenced by the structure or style of regional languages.
-In these cases, we correct the text to adhere to correct Indonesian grammar.
-To maintain the quality of IndoCulture, we rigorously filter instances that display (1) poor writing, as enhancing their quality substantially is difficult; (2) obvious answer options, which allow for easy guessing of the correct choice without understanding the cultural contexts; and (3) ambiguous contexts, where all options are equally valid as the correct answer.
-For example, in a topic about breakfast, the three options might include one traditional food alongside two other very commonly consumed foods in Indonesia.
-Furthermore, we manually verify the province-specific annotation for each instance using the Google search engine.
-We annotate whether the instance pertains to national-level culture.
+In stage 1, we implemented quality control by pairing two annotators from the same province.
+Each annotator was tasked with answering a set of questions prepared by the other annotator, and vice versa.
+During this phase, the annotator were presented with a premise sentence and three shuffled options.
+They were allowed to search for the answer from any source if they were unsure.
+Instances that were incorrectly answered by the second annotator were discarded, as we hypothesize that these instances may contain incorrect answers or possess a level of ambiguity.
+Additionally, annotators were required to identify whether the instance is province-specific (binary annotation: True/False), indicating that it is uniquely relevant in their province and not in others.
+In stage 2 of quality control, the first two authors of this paper performed post-editing of data that passed the first stage of quality control.
+We first focused on correcting the linguistic aspects of the text, including checking for spelling errors.
+Although the text is written in Indonesian, some annotators may use dialects or be influenced by the structure or style of regional languages.
+In these cases, we corrected the text to adhere to Indonesian grammar.
+To maintain the quality of IndoCulture, we rigorously filtered instances that contained: (1) poor writing, in the case that it was difficult to post-edit to enhance their quality; (2) obvious answer options, which allow for easy guessing of the correct choice without understanding the cultural context; and (3) ambiguous contexts, where all options are equally valid as the correct answer.
+For example, in a topic about breakfast, the three options might include one traditional food alongside two other very commonly consumed foods in Indonesia, and be considered too obvious.
+Furthermore, we manually verified the province-specific annotations for each instance using the Google search engine.
+We annotated whether the instance pertains to national-level culture or not.
 If the example is specific to a province, it will be annotated as uncommon in national culture, and vice versa.
 3.2 Data Statistics
-After the annotation process, we initially collected 3,162 instances.
-[^4] Although we requested each annotator to produce 150 instances, not all were able to complete their allotted tasks within the given timeframe.
+After the instance writing process, we initially collected 3,162 instances out of a target of 3,300 instances (22 workers $\times$ 150 subtopics).
+Although we requested each annotator to produce 150 instances, not all were able to complete their allotted tasks within the given timeframe.
 Unfortunately, we were unable to find additional candidates from the same local province to address the data deficiencies Winata et al.
 (2023).
-In stage 1 of quality control, the initial pool of 3,162 instances was reduced to 2,801 instances.
-Our thorough stage 2 of quality control further reduced the sample to 2,429 final high-quality samples.
+In stage 1 of quality control, the initial pool of 3,162 instances was reduced to 2,801 instances, and stage 2 of quality control further reduced the sample to 2,429 high-quality samples.
 The data distribution of IndoCulture per province is presented in Table 2.
@@
 4.1 Set-Up
-We focus on evaluating 23 language models in zero-shot settings: (1) 15 open-source multilingual language models including BLOOMZ Muennighoff et al.
+We evaluate 27 language models in zero-shot settings: (1) nineteen open-weight multilingual language models of varying sizes, namely BLOOMZ Muennighoff et al.
 (2023), mT0 Muennighoff et al.
-(2022), LlaMA2 Touvron et al.
 (2023), Bactrian-X Li et al.
-(2023a) in different sizes, (2) 2 South East Asian language models including SeaLLM Nguyen et al.
-(2023b), and SeaLion Singapore (2023), (3) 4 Indonesian-centric language models, including IndoBART Cahyawijaya et al.
+(2023), Llama–2 Touvron et al.
+(2023), and Llama–3 Dubey et al.
+(2024); (2) two South East Asian language models, namely SeaLLM Nguyen et al.
+(2024), and SeaLion Singapore (2023); (3) four Indonesian-centric language models, namely IndoBART Cahyawijaya et al.
 (2021), IndoGPT Cahyawijaya et al.
-(2021), Merak Ichsan (2023), Komodo Owen et al.
-(2024), (3) 2 closed-source models, including ChatGPT: gpt-3.5-turbo Ouyang et al.
+(2021), Merak Ichsan (2023), and Komodo Owen et al.
+(2024); (3) two closed-weight models, namely ChatGPT: gpt-3.5-turbo Ouyang et al.
 (2022) and GPT–4: gpt-4-0613 OpenAI (2023).
 Please refer to Appendix A for further details.
-Initially, we evaluate the effectiveness of sentence completion and multiple-choice question strategies in predicting the correct options.
+First, we evaluate the effectiveness of sentence completion and multiple-choice question strategies in predicting the correct options using the Indonesian and English prompt templates shown in Figure 3.
 In both scenarios, we conduct benchmarks across three distinct location contexts.
-Formally, given a premise $s$, three candidate options $c_{1}$, $c_{2}$, $c_{3}$, and location $l\in\{\texttt{none},\texttt{Indonesia},p_{j}\}$, for sentence completion, we select the correct option based on:
+Formally, given a premise $s$, three candidate options $c_{1}$, $c_{2}$, $c_{3}$, and location $l\in\{\texttt{none},\texttt{Indonesia},province\}$, for sentence completion, we select the correct option based on:
 […]
@@
 In the case of multiple-choice questions, we devise a template for the prompt question and determine the answer by selecting the option with the highest probability among letters A, B, and C.
-Figure 3 displays the prompt templates in Indonesian and English for both sentence completion and multiple-choice question.
+For GPT–3.5 and GPT–4, we exclude experiments with sentence completion because the closed-weight models do not provide an overall probability score.
+For multiple-choice questions, we select the first generated token that corresponds to the letters A, B, or C using a regular expression.
 Figure 3: Templates for sentence completion and multiple-choice questions prompts.
@@
 $l=\text{None}$ $l=\text{Ind}$ $l=\text{Prov}$ $l=\text{None}$ $l=\text{Ind}$ $l=\text{Prov}$
-Human 100.0 100.0
+Human 100 100.0
 Random 33.3 33.3 33.3 33.3 33.3 33.3
-BLOOMZ (560M) 37.5 35.3 35.3 32.1 32.0 32.1
-BLOOMZ (1.1B) 36.2 36.9 37.2 32.0 32.0 32.0
-BLOOMZ (3B) 38.5 40.5 41.2 45.9 47.6 48.6
-BLOOMZ (7B) 40.9 43.8 44.4 48.8 49.8 49.7
+BLOOMZ (560M) 37.2 35.3 35.3 32.5 32.4 32.5
+BLOOMZ (1.1B) 36.3 36.9 37.2 32.4 32.4 32.4
+BLOOMZ (3B) 38.6 40.7 41.5 47.0 48.6 49.2
+BLOOMZ (7B) 41.3 44.1 44.6 49.5 50.6 50.5
 […]
@@
 […]
-LLaMA2 (7B) 36.7 37.0 37.2 40.0 39.0 37.7
-LLaMA2-chat (7B) 36.9 36.9 37.4 40.5 41.1 40.3
-LLaMA2 (13B) 39.4 39.7 39.9 46.9 46.9 46.6
-LLaMA2-chat (13B) 38.2 38.4 38.9 47.2 49.2 49.2
 […]
 […]
-IndoBART (132M) 41.9 41.1 41.7 32.1 32.0 32.2
-IndoGPT (117M) 42.6 42.1 42.7 33.7 34.1 35.0
-Merak (7B) 40.2 40.6 42.6 51.5 52.5 52.5
-SeaLLM (7B) 39.1 39.7 41.1 51.7 52.4 52.5
-SEA-LION (7B) 38.8 38.7 39.6 33.4 32.8 33.0
+[…]
+Llama–2 chat (7B) 37.3 37.4 37.9 40.6 41.3 40.7
+Llama–2 (13B) 39.6 40.2 40.2 47.6 47.6 47.3
+Llama–2 chat (13B) 38.6 38.9 39.3 47.8 49.6 49.6
+[…]
+Llama–3 Instruct (8B) 41.9 41.5 42.3 56.7 57.6 59.0
+Llama–3 (70B) 51.2 51.7 54.3 68.6 69.9 72.7
+Llama–3 Instruct (70B) 49.2 49.6 52.2 68.5 69.3 73.3
+IndoBART (132M) 42.4 41.3 42.1 32.6 32.4 32.7
+IndoGPT (117M) 42.6 41.9 42.4 33.7 33.8 34.7
+Merak (7B) 41.0 41.5 43.5 51.9 53.1 53.2
+SeaLLM (7B) 39.1 39.3 41.1 52.2 53.1 53.0
+SEA-LION (7B) 38.8 38.9 39.7 33.8 33.0 33.3
 Komodo (7B) 45.1 45.4 46.1 37.6 35.1 36.1
-GPT–3.5 (NA) 58.4 60.2 61.7
-GPT–4 (NA) 68.8 71.5 75.8
+GPT–3.5 (NA) 59.8 60.9 62.7
+GPT–4 (NA) 69.1 71.8 75.9
 Table 3: Zero-shot accuracy across various models and settings.
-“MCQ” refers to the multiple-choice question method, and $l$ denotes the location as additional context.
+“MCQ” refers to the multiple-choice question method, and $l$ denotes the location as additional context (“Ind” and “Prov” denote the country of Indonesia, and the corresponding province).
 The bold numbers highlight the highest score within each model group.
@@
 […]
-Multiple-choice question method is generally better.
+The multiple-choice question method is generally better.
 Our findings suggest that the multiple-choice question method tends to outperform the sentence completion method, with exceptions noted for BLOOMZ (560M, 1.1B), IndoBART, IndoGPT, and Komodo.
-Interestingly, within the sentence completion method, the small Indonesian-centric language model, IndoGPT, surpasses the performance of almost all large multilingual models and remains competitive with BLOOMZ (7B).
-However, IndoGPT’s performance experiences a significant decline in the multiple-choice question method, with a notable margin of nearly 8 to 10 points.
-While Komodo (7B) achieves the highest accuracy with the sentence completion method, its performance also decreases by 10-12 points in the multiple-choice method.
+Interestingly, in the sentence completion task, the Indonesian-focused language model Komodo (7B) outperforms nearly all large multilingual models, with the exception of Llama–3 (70B).
+However, Komodo experiences a significant decline for the multiple-choice question method, with a notable margin of 10–12 points.
 This discrepancy could potentially be attributed to differences in the nature of language model training and instruction-tuning.
 Impact of location context on model performance
-Our investigation reveals that incorporating various levels of location granularity has a noticeable effect on zero-shot performances, especially models with larger parameter sizes.
-Detailed location contexts notably enhance the accuracy of BLOOMZ (7B), LlaMA2 (13B), Merak (7B), SeaLLM (7B), Komodo (7B), GPT–3.5, and GPT–4.
-For instance, in GPT–4, the accuracy gap between $l=\texttt{none}$ and $l=\texttt{Indonesia}$ is 2.7, and this gap further increases by 7 points when $l=\texttt{Prov}$ is assigned.
+Our investigation reveals that incorporating various levels of location granularity has a noticeable effect on zero-shot performance, especially for models with larger parameter sizes.
+Detailed location context notably enhances the accuracy of BLOOMZ (7B), Llama–2 (13B), Llama–3 (70B), Merak (7B), SeaLLM (7B), Komodo (7B), GPT–3.5, and GPT–4.
+For instance, in GPT–4, the accuracy gap between $l=\texttt{none}$ and $l=\texttt{Indonesia}$ is 2.7, and this gap further increases by 7 points when $l=\texttt{Province}$ is assigned.
 4.3 Analysis
-Given the superior performance of models with large parameter sizes using the multiple-choice question method and location $l=\texttt{Prov}$, we employ these settings for analysis.
-Province Merak GPT–3.5 GPT–4
+Given the exceptional performance of models with large parameter sizes using the multiple-choice question method and location $l=\texttt{Province}$, we employ these configurations for our analysis.
+In this section, our main focus is on the top three performing models: Llama–3 Instruct (70B), Merak (7B), and GPT–4.
+These models represent a multilingual open-weight model, Indonesian-centric open-weight model, and closed-weight model, respectively.
+Province Merak Llama–3 GPT–4
 $\neg$¬PS PS $\neg$¬PS PS $\neg$¬PS PS
-Aceh 59.7 53.4 79.2 65.5 93.1 73.6
-North Sumatra 58.3 46.9 63.9 57.2 75.0 73.2
-West Sumatra 50.0 41.3 63.2 48.4 85.5 63.7
-West Java 65.3 58.8 84.2 71.8 92.6 81.7
-Cental Java 58.8 47.1 67.6 52.9 82.4 72.9
-East Java 57.4 37.6 72.2 46.5 87.0 63.4
-Bali 78.9 65.6 87.7 81.7 93.0 89.4
-NTT 64.3 52.0 57.1 49.3 85.7 68.0
-South Borneo 65.7 46.4 68.6 53.6 77.1 67.4
-South Sulawesi 53.3 47.9 66.7 55.8 80.0 70.6
-Papua 76.7 55.2 76.7 62.8 90.0 71.3
+Aceh 59.7 53.4 77.8 68.4 93.1 73.6
+North Sumatra 58.3 46.9 69.4 67.5 75.0 73.2
+West Sumatra 50.0 41.3 73.7 64.6 85.5 63.7
+West Java 65.3 58.8 91.6 86.3 92.6 81.7
+Cental Java 58.8 47.1 79.4 67.1 82.4 72.9
+East Java 57.4 37.6 88.9 57.4 87.0 63.4
+Bali 78.9 65.6 96.5 86.7 93.0 89.4
+NTT 64.3 52.0 78.6 61.3 85.7 68.0
+South Borneo 65.7 46.4 85.7 66.9 77.1 67.4
+South Sulawesi 53.3 47.9 73.3 66.9 80.0 70.6
+Papua 76.7 55.2 90.0 73.1 90.0 71.3
 Table 4: Top-3 model accuracy by province.
-"PS" indicates instances containing province-specific contexts, while "$\neg$¬PS" indicates otherwise.
+“PS” indicates instances containing province-specific context, while “$\neg$¬PS” indicates otherwise.
 The green and red cells indicate the top three and bottom three scores, respectively.
 Results by province
-Table 3 highlights that the top 3 performing LLMs exhibit a nuanced understanding of cultures within Indonesian provinces, particularly excelling in the cultures of West Java and Bali compared to other provinces.
-GPT–4, for instance, achieves the best accuracy of 93%, while Merak and GPT–3.5 also demonstrate strong performance in Bali and West Java, with accuracies ranging between 55% and 88%.
+Table 4 highlights that the top 3 performing LLMs exhibit a nuanced understanding of culture within Indonesian provinces, particularly excelling in the cultures of West Java and Bali compared to other provinces.
+Llama–3 and GPT–4, for instance, achieve the best accuracy of more than 90%, while Merak achieves the best performance in Bali, Papua, and West Java, with accuracies ranging between 55% and 79%.
 In other provinces like West Sumatra and South Borneo, the models typically exhibit poorer performance.
-Specifically, for GPT-3.5, the performance gap compared to Bali ranges from 19 to 33%.
+Specifically, for Llama–3, the performance gap compared to Bali ranges from 10 to 30%.
 This highlights the presence of cultural biases and a lack of inclusivity in model reasoning abilities, likely stemming from the distribution of training data.
@@
 Results by topic
-Table 4 shows the accuracy of the top 3 performing models across topics.
-Similar to Table 3, the models perform better in province-specific context for all topics, with death customs being an exception.
-For province-specific contexts, GPT–4 excels in themes of wedding, death, and arts while for non-specific contexts, GPT–4 achieves accuracy more than 90.0 for family relationship, religious holidays, and farm and fishery.
-Province Merak GPT–3.5 GPT–4
+Table 5 shows the accuracy of the top 3 performing models across different topics.
+Similar to Table 4, the models perform worse in province-specific contexts for all topics, with the notable exception of food.
+For province-specific contexts, GPT–4 excels for the themes of food, religious holidays, and arts while for non-specific contexts, Llama–3 achieves accuracy more than 90% for agriculture, daily activities, and religious holidays.
+Topic Merak Llama–3 GPT–4
 $\neg$¬PS PS $\neg$¬PS PS $\neg$¬PS PS
-Food 56.9 51.0 65.9 62.1 79.3 63.5
-Wedding 66.7 64.3 77.6 56.0 92.8 78.6
-Family relationship 65.5 54.5 72.5 53.9 93.1 72.7
-Pregnancy and kids 67.9 42.9 69.2 50.6 82.1 72.7
-Death 38.5 55.2 60.9 62.5 76.9 82.5
-Religious holidays 68.5 48.8 79.3 68.8 94.4 64.6
-Agriculture 70.5 52.3 71.4 62.1 88.5 76.1
-Farm and fishery 77.8 51.0 77.8 51.9 94.4 75.8
-Art 58.7 50.2 85.2 61.1 78.3 77.3
-Traditional game 60.0 50.2 80.3 59.8 88.0 66.9
-Daily activities 55.1 39.6 66.0 64.8 87.8 63.0
-Socio-religious 57.1 44.3 79.6 61.1 85.7 64.6
+Food 58.7 50.2 69.6 71.8 73.3 77.3
+Wedding 60.0 50.2 80.0 66.9 82.5 66.9
+Family relationship 55.1 39.6 75.5 66.9 85.2 63.0
+Pregnancy and kids 57.1 44.3 83.5 69.0 84.7 64.6
+Death 56.9 51.0 86.2 62.5 79.4 63.5
+Religious holidays 66.7 64.3 79.7 78.6 91.5 78.6
+Agriculture 65.5 54.5 100.0 80.3 93.8 72.7
+Farm and fishery 67.9 42.9 85.7 62.3 80.0 72.7
+Art 77.8 55.2 83.3 72.2 90.0 82.5
+Traditional game 38.5 48.8 69.2 68.3 78.6 64.6
+Daily activities 68.5 52.3 92.6 72.7 93.1 76.1
+Socio-religious 70.5 51.0 88.5 71.8 86.4 75.8
 Table 5: Top-3 model accuracy by topic.
-"PS" indicates instances containing province-specific contexts, while "$\neg$¬PS" indicates otherwise.
+“PS” indicates instances containing province-specific contexts, while “$\neg$¬PS” indicates otherwise.
 The green and red cells indicate the top three and bottom three scores, respectively.
+Results by fine-grained cultural elements
+We tasked two expert workers with annotating 200 random samples from IndoCulture based on six cultural elements, as derived from Axtell and Fornwald (1998); Williams (2014).
+[^3] https://www.languageeducatorsassemble.com/elements-of-culture [^4] https://pressbooks.howardcc.edu/soci101/chapter/3-2-the-elements-of-culture/ While these elements may not encompass every cultural aspect, we contend that they cover the most prominent or pivotal elements, including: (1) symbols (material or non-material objects representing meaning); (2) artifacts (material or non-material objects produced by society); (3) values and beliefs (principles, ideas, and concepts assumed to be ideal and correct in society); (4) norms (rules guiding values and beliefs); (5) language; (6) rituals (established procedures and ceremonies); and (7) other, for examples that do not fit into any of the defined elements.
+This annotation is a multi-label task, and the average Kappa score across the cultural elements is 0.56, with each ranging from 0.4 to 0.75.
+These scores indicate moderate to substantial agreement.
+Table 6 displays the distribution of each cultural element in our dataset, along with the performance breakdown across Merak, Llama–3, and GPT–4.
+Among the 200 random samples, we observe that 42.5% of our data contains artifacts, 37.5% norms, and 30% rituals.
+Only 4% of the data pertains to symbols, while 7.5% belongs to the other category.
+Merak shows lower accuracy in norms, with a 24% decrease compared to values and beliefs.
+Conversely, Llama–3 performs best in values and beliefs with 85% accuracy, but accuracy drops by 23% for norms.
+GPT–4 maintains relatively stable accuracies across cultural elements, with differences averaging between 3% and 5%.
+Furthermore, language presents a challenge for Merak, achieving only 38% accuracy, whereas Llama–3 and GPT–4 achieve 66% and 72% accuracy, respectively.
+Cultural element (%) Merak Llama–3 GPT–4
+Symbols (4) 50.0 50.0 70.8
+Artifacts (42.5) 55.3 74.1 67.8
+Values and Beliefs (10.5) 61.9 85.7 69.8
+Norms (37.5) 38.7 62.7 73.6
+Language (19.5) 38.5 66.7 72.0
+Ritual (30) 53.3 65.0 70.7
+Other (7.5) 66.7 66.7 69.2
+Table 6: Accuracy comparison of Merak, Llama–3–Instruct (70B), and GPT–4 across 200 random samples, categorized by cultural elements.
+The numerical value following each cultural element indicates its proportion within the samples.
+Can the model provide a reasonable explanation to support the answer?
+Figure 4: Performance comparison between Merak (7B), Llama–3 Instruct (70B), and GPT–4 based on text generation output.
+“Answer (T)” indicates that the generated answer is true, while “Exp(F)” denotes that the answer explanation is false.
+We conduct a manual investigation of the text generation output for Merak (7B), Llama–3, and GPT–4 across 200 random samples.
+This involves manually examining the generated answer along with its explanation.
+To obtain the explanation, we modify the Indonesian prompt in Figure 3 by adding the string Jelaskan jawabanmu!
+“Explain your answer!”.
+Our annotation process is binary, categorizing explanations as either True or False.
+We label an explanation as False if it is absent, contains hallucinations, or provides inaccurate information.
+[^5] We use the Google search engine to verify the correctness of the explanation.
+As anticipated, there is a substantial drop in accuracy for Merak (7B) from 53.2% (as shown in Table 3) to 29.5%.
+This discrepancy underscores the limitations of relying solely on token probabilities to assess the true capability of a language model.
+Interestingly, only 4.5% of the samples are answered correctly with the appropriate explanation by Merak, despite it being the top performer among the Indonesian-centric language models.
+Larger models like Llama–3 and GPT–4 achieve more robust accuracies of 70.5% and 69.5%, which are 3–5% lower than those indicated in Table 3.
+However, both models encounter challenges in generating appropriate explanations for correctly-answered samples, with Llama–3 and GPT–4 producing explanation errors in 29% and 34% of cases, respectively.
 Does language affect model performance?
@@
 and used the English prompt in Figure 3 to evaluate the models.
-All model performances in English text dropped except for LLaMA-2 and Merak.
+Specifically for this part, we include more models for comparison.
+All results over English text dropped, except for Llama–2 and Merak.
 This could be attributed to two reasons.
-First, LLaMA-2 is an English-centric model, and Merak is pretrained from LLaMA-2.
-Second, the performance drop in other models could be caused by translation errors.
+First, Llama–2 is an English-centric model, and Merak is fine-tuned from Llama–2.
+Second, the performance drop for other models could be caused by translation errors.
 We further investigated this with 100 random samples and found that 81 samples had acceptable translations.
-We observed translation errors such as pronoun mismatches, inaccurate proverb translations, and inaccurate translations of local terms, such as “pupuik” translated as “fertilizer”.666pupuik is a traditional musical instrument in West Sumatra.
+We observed translation errors such as pronoun mismatches, inaccurate proverb translations, and inaccurate translations of local terms, such as pupuik translated as “fertilizer”.777pupuik is a traditional musical instrument in West Sumatra.
 It is worth noting that the word pupuik closely resembles the word pupuk in Indonesian, which means fertilizer.
 To better understand the cultural gap in language models, we followed the approach of Liu et al.
-(2023a) to manually correct the translations and reevaluate the models.
-We found that GPT-4’s performance over the 100 random samples was 77.0 for the original Indonesian text, 68.0 for the English machine-translated text, and 72.0 for the English translation fixed by humans.
+(2024) to manually correct the translations and reevaluate the models.
+We found that GPT–4’s performance over the 100 random samples was 77.0 for the original Indonesian text, 68.0 for the English machine-translated text, and 72.0 for the English translation fixed by humans.
 […]
-Results by fine-grained culture elements
-We tasked two expert workers with annotating 200 random samples from IndoCulture based on six cultural elements, as derived from Axtell and Fornwald (1998); Williams (2014).
-[^7] https://www.languageeducatorsassemble.com/elements-of-culture [^8] https://pressbooks.howardcc.edu/soci101/chapter/3-2-the-elements-of-culture/ While these elements may not encompass every cultural aspect, we contend that they cover the most prominent or pivotal elements, including: (1) symbols (material or non-material objects representing meaning), (2) artifacts (material or non-material objects produced by society), (3) values and beliefs (principles, ideas, and concepts assumed to be ideal and correct in society), (4) norms (rules guiding values and beliefs), (5) language, (6) rituals (established procedures and ceremonies), and (7) other, for examples that do not fit into any of the defined elements.
-This annotation is a multi-label task, and the average Kappa score across the cultural elements is 0.56, with each ranging from 0.4 to 0.75.
-These scores indicate moderate to substantial agreement.
-Table 5 displays the distribution of each cultural element in our dataset, along with the performance breakdown across Merak, GPT-3.5, and GPT-4.
-Among the 200 random samples, we observe that 42.5% of our data contains artifacts, 37.5% norms, and 30% rituals.
-Only 4% of the data pertains to symbols, while 7.5% belongs to the other category.
-Merak exhibits lower performance in norms, with a 24% decrease compared to values and beliefs.
-Conversely, GPT-3.5 and GPT-4 perform better in norms, with the performance gap in values and beliefs being only $-4$ points.
-Additionally, language poses a challenge in Merak, with only 38% accuracy, while GPT-3.5 and GPT-4 achieve 60% and 72% accuracy, respectively.
-Cultural element (%) Merak GPT–3.5 GPT-4
-Symbols (4) 50.0 58.9 70.8
-Artefacts (42.5) 55.3 59.1 67.8
-Values and Beliefs (10.5) 61.9 57.5 69.8
-Norms (37.5) 38.7 63.2 73.6
-Language (19.5) 38.5 60.2 72.0
-Ritual (30) 53.3 60.0 70.7
-Other (7.5) 66.7 58.9 69.2
-Table 5: Accuracy comparison of Merak, GPT–3.5, and GPT–4 across 200 random samples, categorized by cultural elements.
-The numerical value following each cultural element indicates its proportion within the samples.
-Can the model provide a reasonable explanation to support the answer?
-Figure 5: Performance comparison between Merak (7B) and GPT–4 based on text generation output.
-“Answer (T)” indicates that the generated answer is true, while ”Exp(F)” denotes that the answer explanation is false.
-We conduct a manual investigation of the text generation output for Merak (7B) and GPT–4 across 200 random samples.
-This involves manually examining the generated answer along with its explanation.
-To obtain explanation, we modify the Indonesian prompt in Figure 3 by adding the string Jelaskan jawabanmu!
-("Explain your answer!").
-Our annotation process is binary, categorizing explanations as either True or False.
-We label an explanation as False if it is absent, contains hallucinations, or provides inaccurate information.
-[^9] We use Google search engine to verify the correctness of the explanation.
-As anticipated, there is a significant drop in accuracy for Merak (7B) from 52.5% (as shown in Table 2) to 29.5%.
-This discrepancy underscores the limitations of relying solely on token probability to assess the true capability of a language model.
-Interestingly, only 4.5% of the samples are answered correctly with the appropriate explanation by Merak, despite it being the top performer among the open-source models.
-GPT–4 achieves an accuracy of 69.5%, which is 5 points lower than indicated in Table 2, possibly due to a smaller sample size and biases in regular expressions.
-However, GPT–4 provides better explanations, with 51.5% of the samples correctly answered along with the appropriate explanation.
 5 Discussion
-The recent study Wang et al.
-(2024) has demonstrated that evaluating language models using token probability in multiple-choice question types does not align well with the generated text.
+A recent study Wang et al.
+(2024) has demonstrated that evaluating language models using token probabilities in multiple-choice question types does not align well with the generated text.
 This discrepancy is reported to be more pronounced in models fine-tuned on conversational or safety data.
-In response to this issue, we have conducted a manual evaluation on 200 random samples (in Section 4.3) and found that the performance in generated texts, especially for Merak, the best open-source language model, is deteriorating.
-However, this issue is less apparent in GPT-4, as the initial evaluation is based on regular expressions.
+In response to this issue, we conducted a manual evaluation of 200 random samples (in Section 4.3) and found that the performance in the generated text, especially for Merak — the best open-weight Indonesian-centric language model — deteriorates.
+However, this issue is less apparent in larger models such as Llama–3 and GPT–4.
 Conducting manual evaluation on all data and models is expensive, and we plan to address this issue in future works.
-This work primarily focuses on introducing a novel dataset meticulously constructed for evaluating cultural commonsense reasoning within the Indonesian context, including preliminary evaluation results based on the standard methods used in previous studies OpenAI (2023); Touvron et al.
+This work primarily focuses on introducing a novel dataset constructed for evaluating cultural commonsense reasoning within the Indonesian context, including preliminary evaluation results based on standard methods used in previous studies OpenAI (2023); Touvron et al.
 (2023); Koto et al.
 (2023); Li et al.
-(2023b); Koto et al.
+(2024); Koto et al.
 (2024).
@@
 Through collaboration with local individuals across eleven provinces and rigorous quality control measures, we introduce IndoCulture for the purpose of evaluating language models.
-Our findings reveal that large language models, whether Indonesian-centric or multilingual, demonstrate a limited understanding of Indonesian culture.
-Notably, incorporating location as additional context significantly enhances model performance, particularly in GPT–3.5 and GPT–4.
+Our findings reveal that large language models, whether Indonesian-centric or multilingual, demonstrate a limited understanding of Indonesian cultures.
+Notably, incorporating location as additional context significantly enhances model performance, particularly for GPT–3.5 and GPT–4.
 Limitations
@@
 It does not consider temporal aspects.
-Our dataset was created in the year 2023, recognizing that cultures may evolve over time, as discussed by Mesoudi (2016).
+Our dataset was created in the year 2023, and we recognize that cultures may evolve over time, as discussed by Mesoudi (2016).
 Furthermore, as demonstrated in Section 4.3, a significant portion of IndoCulture comprises cultural elements such as artifacts, norms, and rituals.
@@
 We encourage future research to further explore these cultural elements and to expand the geographical coverage beyond the eleven provinces studied in this paper.
+We also acknowledge that our dataset size is relatively small, to cover all the provinces.
+However, compared to existing Indonesian datasets that have been manually curated by natives (see COPAL–ID in Table 1), ours is significantly larger in terms of both size and regional coverage.
+Future work may extend the size and the coverage of IndoCulture to get a more holistic picture of Indonesian cultures.
 Ethical Considerations
@@
 [^8] https://creativecommons.org/licenses/by-nc-sa/4.0/ Our data is intended for academic research and non-commercial purposes.
-Workers are compensated above the minimum monthly salary in Indonesia and are fully aware that the data will be released to the public.
+Workers were compensated above the minimum monthly salary in Indonesia and are fully aware that the data will be released to the public.
 It is important to note that no private or sensitive information of the workers is included in IndoCulture.
-Appendix A Model Artifacts
+Acknowledgements
+We acknowledge the thorough feedback and impactful suggestions from the reviewers and the action editor.
+We also extend our gratitude to the 22 annotators from eleven provinces in Indonesia for their valuable contributions in constructing IndoCulture.
+This project was supported by MBZUAI and UI through PUTI-Q2 grant (NKB-1192/UN2.RST/HKP.05.00/2022).
+Appendix A Model Details
 Models (#parameters) Source
@@
 […]
-LLamA2-chat (7B) meta-llama/Llama-2-7b-chat
-LLamA2 (13B) meta-llama/Llama-2-13b
-LLamA2-chat (13B) meta-llama/Llama-2-13b-chat
+Llama–2 chat (7B) meta-llama/Llama--2-7b-chat
+Llama–2 (13B) meta-llama/Llama--2-13b
+Llama–2 chat (13B) meta-llama/Llama--2-13b-chat
+[…]
+Llama–3 Instruct (8B) meta-llama/Meta-Llama--3-8B-Instruct
+Llama–3 (70B) meta-llama/Meta-Llama--3-70B
+Llama–3-chat (70B) meta-llama/Meta-Llama--3-70B-Instruct
 […]
@@
 Komodo (7B) Yellow-AI-NLP/komodo-7b-base
-Table 6: With the exception of GPT-3.5 and GPT-4, all the models used in this study were sourced from Huggingface Wolf et al. (2020).
+Table 7: With the exception of GPT–3.5 and GPT–4, all the models used in this study were sourced from Huggingface Wolf et al. (2020).